OpenStudy (anonymous):

let f(x)=sqrt[3]{x-5}. find f^-1(x).

OpenStudy (anonymous):

y= mx+b

OpenStudy (anonymous):

the answer would be f^-1(x)=(x+5)^3, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

Just making sure. Thanks!

OpenStudy (zarkon):

no

OpenStudy (anonymous):

no?

OpenStudy (agreene):

nope

OpenStudy (asnaseer):

assuming you don;t need to specify the valid domain of x

OpenStudy (anonymous):

hold on

OpenStudy (asnaseer):

sorry - it shud be x^3+5

OpenStudy (asnaseer):

I mis-read your answer

OpenStudy (zarkon):

if \[f(x)=\sqrt[3]{x-5}\] then \[f^{-1}(x)=x^3+5\]

OpenStudy (anonymous):

wait... hm...

OpenStudy (anonymous):

i think it really depends on how you read it

OpenStudy (anonymous):

these are the possible answers: a) f^-1(x)=3x-15 b) f^-1(x)=(x+5)^3 c) f^-1(x)=x^3+5 d) f^-1(x)=5/3 e) f(x) is not one-to-one, therefore it does not have an inverse

OpenStudy (agreene):

c.

OpenStudy (zarkon):

there is no other way to read it

OpenStudy (asnaseer):

@Zarkon and @agreene are right

OpenStudy (anonymous):

ok could explain why C is the answer

OpenStudy (zarkon):

reported

OpenStudy (asnaseer):

reported

OpenStudy (zarkon):

mods can see deleted posts

OpenStudy (asnaseer):

sorry for the confusion @mariomintchev - I should not have rushed to the answer initially.

OpenStudy (zarkon):

the point is that just because you delete a post doesn't mean it is gone forever

OpenStudy (anonymous):

thats fine. just need an explanation.

OpenStudy (anonymous):

Hmmm...

OpenStudy (agreene):

\[y=\sqrt[3]{x-5}\] \[y=(x-5)^{1/3}\] \[y^3=((x-5)^{1/3})^3\] \[y^3=x-5\] \[y^3+5=x\] \[f^{-1}=x^3+5\]

OpenStudy (anonymous):

It’s C.

OpenStudy (anonymous):

right? is it c?

OpenStudy (asnaseer):

@MarkSolves has explained it for you

OpenStudy (anonymous):

mark. we already have the answer. -_-

OpenStudy (anonymous):

ok well sorrreeeeee

OpenStudy (agreene):

I worked it ^^ go look, it's fantastically uninteresting!

OpenStudy (anonymous):

why are you all here if you already have the answer -.-

OpenStudy (asnaseer):

sorry I meant @agreene has explained it - I really must get my glasses cleaned! :-)

OpenStudy (zarkon):

it's a party BYOB though

OpenStudy (agreene):

BYOB? man... that means I need to run to the store >.<

OpenStudy (zarkon):

I have plenty of wine in the house :)

OpenStudy (agreene):

but then its BYOW party... and no one like those.

OpenStudy (asnaseer):

ok folks - lets all get down to @Zarkons!

OpenStudy (anonymous):

@agreene THANKS!

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

@everyone else LOL

OpenStudy (anonymous):

lols

OpenStudy (anonymous):

jk jk!

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

what happened to the comments?